# D³-Predictor: Noise-Free Deterministic Diffusion for Dense Prediction

### v3

@@ 8.4 Generalization to Domain-Specific Tasks
Table 10: Evaluation of the Pavement Crack Detect task [zhang2016road].
@@
SAM [kirillov2023segment] 1.1B 0.351 0.453 0.400
CLIPSeg [luddecke2022image] 345K 0.487 0.499 0.493
Grounded-SAM [ren2024grounded] 1.1B 0.393 0.429 0.428
@@
Table 11: Evaluation of the Auto. Pothole Detect task [fan2019pothole].
@@
SAM [kirillov2023segment] 1.1B 0.162 0.415 0.186
CLIPSeg [luddecke2022image] 345K 0.597 0.819 0.668
Grounded-SAM [ren2024grounded] 1.1B 0.429 0.469 0.460
@@
Table 12: Evaluation of the Pedestrian Safety Monitor task [wang2007object].
@@
SAM [kirillov2023segment] 1.1B 0.417 0.495 0.478
CLIPSeg [luddecke2022image] 345K 0.616 0.717 0.687
Grounded-SAM [ren2024grounded] 1.1B 0.423 0.505 0.478
@@
Table 13: Evaluation of the Precise Road Model task [mnih2013machine].
@@
SAM [kirillov2023segment] 1.1B 0.434 0.465 0.464
CLIPSeg [luddecke2022image] 345K 0.439 0.562 0.505
Grounded-SAM [ren2024grounded] 1.1B 0.278 0.472 0.357
@@
Table 14: Evaluation of the Retinal Vessel Analysis task [jin2022fives].
@@
SAM [kirillov2023segment] 1.1B 0.273 0.594 0.370
CLIPSeg [luddecke2022image] 345K 0.472 0.510 0.499
Grounded-SAM [ren2024grounded] 1.1B 0.344 0.376 0.409
@@
Table 15: Evaluation of the Spinal Morphology Assessment task [chu2015annotated].
@@
SAM [kirillov2023segment] 1.1B 0.555 0.611 0.612
CLIPSeg [luddecke2022image] 345K 0.519 0.569 0.551
Grounded-SAM [ren2024grounded] 1.1B 0.464 0.481 0.481
@@
Table 16: Evaluation of the Oocyte Detect task [letort2022interpretable].
@@
SAM [kirillov2023segment] 1.1B 0.952 0.974 0.969
CLIPSeg [luddecke2022image] 345K 0.366 0.500 0.422
Grounded-SAM [ren2024grounded] 1.1B 0.390 0.518 0.460
@@
Table 17: Evaluation of the Nucleus Localization task [naylor2018segmentation].
@@
SAM [kirillov2023segment] 1.1B 0.277 0.498 0.336
CLIPSeg [luddecke2022image] 345K 0.452 0.511 0.488
Grounded-SAM [ren2024grounded] 1.1B 0.447 0.506 0.496
@@
Table 18: Evaluation of the Water Body Map task [2020SAR].
@@
SAM [kirillov2023segment] 1.1B 0.299 0.415 0.348
CLIPSeg [luddecke2022image] 345K 0.172 0.504 0.235
Grounded-SAM [ren2024grounded] 1.1B 0.322 0.435 0.406
@@
Table 19: Evaluation of the Auto. Tree Tagging task [Tree-Binary-Segmentation].
@@
SAM [kirillov2023segment] 1.1B 0.765 0.840 0.809
CLIPSeg [luddecke2022image] 345K 0.463 0.722 0.579
Grounded-SAM [ren2024grounded] 1.1B 0.391 0.529 0.457
@@
Table 20: Evaluation of the Camouflage Detect task [fan2020camouflaged].
@@
SAM [kirillov2023segment] 1.1B 0.548 0.612 0.601
CLIPSeg [luddecke2022image] 345K 0.569 0.753 0.667
Grounded-SAM [ren2024grounded] 1.1B 0.454 0.504 0.497
@@
Table 21: Evaluation of the Oil Spill Track task [zhu2021oil].
@@
SAM [kirillov2023segment] 1.1B 0.133 0.437 0.187
CLIPSeg [luddecke2022image] 345K 0.427 0.618 0.511
Grounded-SAM [ren2024grounded] 1.1B 0.376 0.490 0.463
@@
Table 22: Evaluation of the Urban Layout Analysis task [mnih2013machine].
@@
SAM [kirillov2023segment] 1.1B 0.386 0.475 0.437
CLIPSeg [luddecke2022image] 345K 0.315 0.616 0.421
Grounded-SAM [ren2024grounded] 1.1B 0.284 0.555 0.390
@@
Table 23: Evaluation of the Weed Distribution Mapping task [haug15].
@@
SAM [kirillov2023segment] 1.1B 0.364 0.399 0.398
CLIPSeg [luddecke2022image] 345K 0.609 0.858 0.725
Grounded-SAM [ren2024grounded] 1.1B 0.126 0.420 0.173
@@
Table 24: Evaluation of the Mask-Wear Monitor task [rezvani2024abanet].
@@
SAM [kirillov2023segment] 1.1B 0.540 0.614 0.584
CLIPSeg [luddecke2022image] 345K 0.377 0.513 0.469
Grounded-SAM [ren2024grounded] 1.1B 0.377 0.498 0.405
@@
Table 25: Evaluation of the Cigarette Detect task [cigarettebutts2024].
@@
SAM [kirillov2023segment] 1.1B 0.304 0.308 0.324
CLIPSeg [luddecke2022image] 345K 0.495 0.500 0.497
Grounded-SAM [ren2024grounded] 1.1B 0.479 0.483 0.486
@@
Table 26: Evaluation of the Fire Alert task [chino2015bowfire].
@@
SAM [kirillov2023segment] 1.1B 0.528 0.591 0.574
CLIPSeg [luddecke2022image] 345K 0.642 0.753 0.710
Grounded-SAM [ren2024grounded] 1.1B 0.469 0.530 0.516
@@
Table 27: Evaluation of the Food Cell Inspect task [letort2022interpretable].
@@
SAM [kirillov2023segment] 1.1B 0.936 0.961 0.959
CLIPSeg [luddecke2022image] 345K 0.291 0.500 0.367
Grounded-SAM [ren2024grounded] 1.1B 0.402 0.561 0.487
@@
Table 28: Evaluation of the Close-Range Segment task [qin2022highly].
@@
SAM [kirillov2023segment] 1.1B 0.412 0.478 0.474
CLIPSeg [luddecke2022image] 345K 0.429 0.498 0.459
Grounded-SAM [ren2024grounded] 1.1B 0.436 0.566 0.516
@@
Table 29: Evaluation of the Cardiac Screen task [cardiacsegmentation2020].
@@
SAM [kirillov2023segment] 1.1B 0.535 0.611 0.612
CLIPSeg [luddecke2022image] 345K 0.519 0.569 0.551
Grounded-SAM [ren2024grounded] 1.1B 0.464 0.481 0.481

### v4

@@ 8.4 Generalization to Domain-Specific Tasks
Table 10: Evaluation of the Pavement Crack Detect task [89].
@@
SAM [35] 1.1B 0.351 0.453 0.400
CLIPSeg [48] 345K 0.487 0.499 0.493
Grounded-SAM [62] 1.1B 0.393 0.429 0.428
@@
Table 11: Evaluation of the Auto. Pothole Detect task [16].
@@
SAM [35] 1.1B 0.162 0.415 0.186
CLIPSeg [48] 345K 0.597 0.819 0.668
Grounded-SAM [62] 1.1B 0.429 0.469 0.460
@@
Table 12: Evaluation of the Pedestrian Safety Monitor task [74].
@@
SAM [35] 1.1B 0.417 0.495 0.478
CLIPSeg [48] 345K 0.616 0.717 0.687
Grounded-SAM [62] 1.1B 0.423 0.505 0.478
@@
Table 13: Evaluation of the Precise Road Model task [52].
@@
SAM [35] 1.1B 0.434 0.465 0.464
CLIPSeg [48] 345K 0.439 0.562 0.505
Grounded-SAM [62] 1.1B 0.278 0.472 0.357
@@
Table 14: Evaluation of the Retinal Vessel Analysis task [29].
@@
SAM [35] 1.1B 0.273 0.594 0.370
CLIPSeg [48] 345K 0.472 0.510 0.499
Grounded-SAM [62] 1.1B 0.344 0.376 0.409
@@
Table 15: Evaluation of the Spinal Morphology Assessment task [12].
@@
SAM [35] 1.1B 0.555 0.611 0.612
CLIPSeg [48] 345K 0.519 0.569 0.551
Grounded-SAM [62] 1.1B 0.464 0.481 0.481
@@
Table 16: Evaluation of the Oocyte Detect task [42].
@@
SAM [35] 1.1B 0.952 0.974 0.969
CLIPSeg [48] 345K 0.366 0.500 0.422
Grounded-SAM [62] 1.1B 0.390 0.518 0.460
@@
Table 17: Evaluation of the Nucleus Localization task [53].
@@
SAM [35] 1.1B 0.277 0.498 0.336
CLIPSeg [48] 345K 0.452 0.511 0.488
Grounded-SAM [62] 1.1B 0.447 0.506 0.496
@@
Table 18: Evaluation of the Water Body Map task [54].
@@
SAM [35] 1.1B 0.299 0.415 0.348
CLIPSeg [48] 345K 0.172 0.504 0.235
Grounded-SAM [62] 1.1B 0.322 0.435 0.406
@@
Table 19: Evaluation of the Auto. Tree Tagging task [39].
@@
SAM [35] 1.1B 0.765 0.840 0.809
CLIPSeg [48] 345K 0.463 0.722 0.579
Grounded-SAM [62] 1.1B 0.391 0.529 0.457
@@
Table 20: Evaluation of the Camouflage Detect task [15].
@@
SAM [35] 1.1B 0.548 0.612 0.601
CLIPSeg [48] 345K 0.569 0.753 0.667
Grounded-SAM [62] 1.1B 0.454 0.504 0.497
@@
Table 21: Evaluation of the Oil Spill Track task [93].
@@
SAM [35] 1.1B 0.133 0.437 0.187
CLIPSeg [48] 345K 0.427 0.618 0.511
Grounded-SAM [62] 1.1B 0.376 0.490 0.463
@@
Table 22: Evaluation of the Urban Layout Analysis task [52].
@@
SAM [35] 1.1B 0.386 0.475 0.437
CLIPSeg [48] 345K 0.315 0.616 0.421
Grounded-SAM [62] 1.1B 0.284 0.555 0.390
@@
Table 23: Evaluation of the Weed Distribution Mapping task [22].
@@
SAM [35] 1.1B 0.364 0.399 0.398
CLIPSeg [48] 345K 0.609 0.858 0.725
Grounded-SAM [62] 1.1B 0.126 0.420 0.173
@@
Table 24: Evaluation of the Mask-Wear Monitor task [63].
@@
SAM [35] 1.1B 0.540 0.614 0.584
CLIPSeg [48] 345K 0.377 0.513 0.469
Grounded-SAM [62] 1.1B 0.377 0.498 0.405
@@
Table 25: Evaluation of the Cigarette Detect task [27].
@@
SAM [35] 1.1B 0.304 0.308 0.324
CLIPSeg [48] 345K 0.495 0.500 0.497
Grounded-SAM [62] 1.1B 0.479 0.483 0.486
@@
Table 26: Evaluation of the Fire Alert task [11].
@@
SAM [35] 1.1B 0.528 0.591 0.574
CLIPSeg [48] 345K 0.642 0.753 0.710
Grounded-SAM [62] 1.1B 0.469 0.530 0.516
@@
Table 27: Evaluation of the Food Cell Inspect task [42].
@@
SAM [35] 1.1B 0.936 0.961 0.959
CLIPSeg [48] 345K 0.291 0.500 0.367
Grounded-SAM [62] 1.1B 0.402 0.561 0.487
@@
Table 28: Evaluation of the Close-Range Segment task [59].
@@
SAM [35] 1.1B 0.412 0.478 0.474
CLIPSeg [48] 345K 0.429 0.498 0.459
Grounded-SAM [62] 1.1B 0.436 0.566 0.516
@@
Table 29: Evaluation of the Cardiac Screen task [34].
@@
SAM [35] 1.1B 0.535 0.611 0.612
CLIPSeg [48] 345K 0.519 0.569 0.551
Grounded-SAM [62] 1.1B 0.464 0.481 0.481